OpenStudy (anonymous):

simplifying (15x^2 + 11x - 14) divided by (5x + 7)

OpenStudy (whpalmer4):

\[\frac{15x^2+11x-14}{5x+7}\]

OpenStudy (anonymous):

3x-2 is the answer

OpenStudy (anonymous):

\[\frac{ 15x ^{2}+21x-10x-14 }{ 5x+7 }=\frac{ 3x(5x+7)-2(5x+7) }{ 5x+7 }=3x-2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)